please be clear with the steps taken and understandable 1. Prove that if f(n) = Θ(n2) for all f(n), then ΣΑ(n)-6(n3). i=1 2. Prove that if f.(n) are linear functions - i.e., that f(
Answer #1

1.

Since fi(n) = Θ(n2), this means

0 <= c1.n2 <= fi(n) <= c2.n2 where c1 and c2 are positive constants.

Multiplying this equation with n, we get

0 <= c1.n2 * n <= n * fi(n) <= c2.n2 * n

or, 0 <= c1.n3 <= >fi(n) <= c2.n3

Hence, >fi(n) = Θ(n3) (proved)

2.

Given that fi(n) = Θ(n), this means that

0 <= c1*n <= fi(n) <= c2*n where c1 and c2 are positive constants.

Let fi(i) = c*n where c1<= c <= c2

Now, ∑ i*fi(n) = (1 + 2 + 3 + ..... + n) * c * n = c * n * (n2 + n)/2 = c * (n3/2 + n2/2)

Clearly, for c1 = c and c2 = 2c and n >= 0, we have

0 <= c1*n3 <= c * (n3/2 + n2/2) <= c3 * n3

Hence, ∑ i*fi(n = Θ(n3)